# Multimodal large language models and physics visual tasks: comparative analysis of performance and costs

### v1

@@ 4 Discussion, limitations and future work
The analysis of performance shows that the best-performing MLLMs are already outperforming post-instruction university student averages on the four tested concept inventories. Some of them are coming close to expert-like performance on some of the inventories. The top-tier models from OpenAI and Google are now exceeding 75% accuracy on our benchmark, suggesting real potential for educational deployment. However, the non-uniform performance across the four underlying concept inventories suggests that even the best models still struggle with many items. While the tested models collectively performed quite well on BEMA and TUG-K, their scores dropped significantly on QMVI and, most notably, FTGOT. These findings suggest a need for future research focused on analyzing how specific visual formats and task types impact model performance, as well as the need for qualitative studies examining the reasoning behind MLLM-selected answers.
The cost analysis reveals that performance does not scale linearly with costs. While some of the most capable models remain relatively cost-effective (e.g., o3, o4 mini), others—like Claude Opus 4—underperformed despite high costs. This suggests that institutions cannot rely solely on pricing tiers (cost per token) or provider reputation as proxies for quality across different educational settings. Conversely, certain mid-range or low-cost models offer a compelling balance of performance and affordability. In particular, Gemini 2.5 Flash with “disabled reasoning” achieved over 60% average accuracy at a fraction of the cost of more expensive models. Our results thus indicate that more affordable models can reach performance levels that are close to those of the best performing models. This has important implications for deploying AI in schools or universities operating under financial constraints. However, freely available or open-weight models (e.g., Gemma 3 series) currently perform well below acceptable thresholds for educational use on physics tasks containing images. These models may be interesting for other roles, but are not yet suitable for student-facing educational applications or assistance with assessment or grading when physics images are involved.
@@
First, our analysis focused exclusively on multiple-choice items from well-established concept inventories. While this design allows for standardization and comparability, it does not capture how MLLMs perform on more open-ended or ill-structured physics tasks, including derivations, written explanations, or lab-based data analysis. Future research should expand to include these formats, which are common in authentic classroom and assessment settings.
Second, our evaluation was quantitative in nature. We scored responses based solely on the selected answer choices, without analyzing the correctness or coherence of the generated output text. This leaves open the possibility that models selected correct answers for the wrong reasons—or, conversely, generated valid reasoning but selected incorrect options. A qualitative investigation of model-generated explanations would be a valuable next step to better understand reasoning quality and error patterns. Such analyses could also help determine which types of visual tasks pose the greatest challenges, and why certain inventories consistently produce lower scores even for otherwise high-performing models. This can be done by collecting new data or re-analyzing the data collected in this project [65].
Third, all the models were evaluated using a static, minimal prompt. We intentionally avoided further prompt engineering to reflect default user scenarios and preserve comparability. However, this likely underestimates the full potential of certain models, particularly those responsive to CoT prompting or domain-specific scaffolding. Future studies could explore the impact of tailored prompts.

### v2

@@ 4 Discussion, limitations and future work
The analysis of performance shows that the best-performing MLLMs are already outperforming post-instruction university student averages on the four tested concept inventories. Some of them are coming close to expert-like performance on some of the inventories. The top-tier models from OpenAI and Google are now exceeding 80% and 75% accuracy, respectively, on our benchmark, suggesting real potential for educational deployment. However, the non-uniform performance across the four underlying concept inventories suggests that even the best models still struggle with many items. While the tested models collectively performed quite well on BEMA and TUG-K, their scores dropped significantly on QMVI and, most notably, FTGOT. These findings suggest a need for future research focused on analyzing how specific visual formats and task types impact model performance, as well as the need for qualitative studies examining the reasoning behind MLLM-selected answers.
The cost analysis reveals that performance does not scale linearly with costs. While some of the most capable models remain relatively cost-effective (e.g., GPT-5 mini, o4 mini and o3 mini), others—like Claude Opus 4—underperformed despite high costs. This suggests that institutions cannot rely solely on pricing tiers (cost per token) or provider reputation as proxies for quality across different educational settings. Conversely, certain mid-range or low-cost models offer a compelling balance of performance and affordability. In particular, GPT-5 mini stands out as a well-performing and low-cost model. GPT-5 nano and Gemini 2.5 Flash with “disabled reasoning” also achieved over 60% average accuracy at a fraction of the cost of more expensive models. Our results thus indicate that more affordable models can reach performance levels that are close to those of the best performing models. This has important implications for deploying AI in schools or universities operating under financial constraints. However, freely available or open-weight models (e.g., Gemma 3 series) currently perform well below acceptable thresholds for educational use on physics tasks containing images. These models may be interesting for other roles, but are not yet suitable for student-facing educational applications or assistance with assessment or grading when physics images are involved.
@@
First, our analysis focused exclusively on multiple-choice items from well-established concept inventories. While this design allows for standardization and comparability, it does not capture how MLLMs perform on more open-ended or ill-structured physics tasks, including derivations, written explanations, or lab-based data analysis that better represent authentic assessment practices in physics classrooms. Future research should expand to include these formats, which are common in authentic classroom and assessment settings.
Second, our evaluation was quantitative in nature. We scored responses based solely on the selected answer choices, without analyzing the correctness or coherence of the generated output text. This leaves open the possibility that models selected correct answers for the wrong reasons—or, conversely, generated valid reasoning but selected incorrect options [33, 37]. A qualitative investigation of model-generated explanations would be a valuable next step to better understand reasoning quality and error patterns. Some models (e.g. OpenAI’s reasoning models) offer so-called “reasoning summaries” as an optional part of their output [76]. A reasoning summary is a condensed recap of the content of the reasoning that the model did in the background, as it employed the internal CoT process hidden from the user, to solve the task at hand. Future research could qualitatively investigate these “reasoning summaries” to get better insights into the strengths and weaknesses of these models. Such analyses could also help determine which types of visual tasks pose the greatest challenges, and why certain inventories consistently produce lower scores even for otherwise high-performing models. For example, follow-up studies could attempt to disentangle model strengths and weaknesses in spatial, symbolic, and graphical reasoning. This can be done by collecting new data, or re-analyzing the data collected in this project [71].
Third, all the models were evaluated using a static, minimal prompt. We intentionally avoided further prompt engineering to reflect default user scenarios and preserve comparability. However, this likely underestimates the full potential of certain models, particularly those responsive to CoT prompting or domain-specific scaffolding. Future studies could explore the impact of tailored prompts. It is important to note, though, that OpenAI explicitly advises against CoT prompting for its reasoning models (e.g., GPT-5 series, o3, o4-mini), stating that the approach is likely to be ineffective or can even worsen the models’ performance [77]. For “non-reasoning” models, on the other hand, performance might benefit from more specialised CoT prompting. However, this would likely also lead to an increased number of generated tokens and consequently a higher cost. Systematic evaluation is needed to determine if CoT prompting with non-reasoning models is economically preferable to running reasoning models. Another consideration that can be relevant from an educational perspective is the transparency of the model output. Using models that do not hide part of their output from the user can be preferred, especially when the focus lies on the process of getting to a solution, not just the final answer. In this respect, educators need to experiment with different models and prompting approaches to achieve the desired behaviour for their praticular use case (e.g., tutoring, grading, feedback).